The motorcyclist crosses a horizontal turn with radius R = 25 m at a speed of V = 50 km / h, being on the slip threshold. Calculate the coefficient of sliding friction and the angle of stability.


1
Expert's answer
2021-01-19T07:10:29-0500

a) In order to motorcycle not to slip the force of friction should provide the necessary centripetal acceleration:


"F_{fr}=F_c,""\mu mg=\dfrac{mv^2}{R},""\mu=\dfrac{v^2}{gR},""\mu=\dfrac{(50\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s})^2}{9.8\ \dfrac{m}{s^2}\cdot 25\ m}=0.78"

b) We can find the angle of stability (the angle at which the motorcyclist doesn't slip) from the formula:


"tan\theta=\mu,""\theta=tan^{-1}\mu=tan^{-1}(0.78)=38^{\circ}."

Answer:

a) "\mu=0.78"

b) "\theta=38^{\circ}."